Write a slope-intercept equation for a line parallel to the line x – 5y = 2 which passes through the point (15, –2)
Answer by Alan3354(69443)

Parallel lines have the same slope.
---
Use y-y1 = m*(x-x1) where m = slope and (x1,y1) is the point.